# Machine learning-based identification of Gaia astrometric exoplanet orbits

### v1

@@ 4.2 Exoplanet candidates
The companion of HD 40503 (Gaia-ASOI-002, Gaia Collaboration et al., 2023b; Holl et al., 2023; Marcussen & Albrecht, 2023; Winn, 2022) continues to await full RV confirmation. The companion of * 54 Cas (Gaia-ASOI-003) is an exoplanet candidate discussed in Gaia Collaboration et al. (2023b). The star BD$+$75 510 (HIP 66071, Gaia-ASOI-001, Holl et al., 2023; Gaia Collaboration et al., 2023b) was followed up with high-precision RVs (Sozzetti et al., 2023) and a sub-Jupiter mass exoplanet was discovered. We also identify HD 128717 (Gaia-ASOI-009) as a strong candidate for hosting a super-Jupiter, motivating prioritised follow-up of this star.
The companion of BD$+$24 4592 (Gaia-ASOI-005) appears to be a strong exoplanet candidate in our analysis on the basis of the Gaia DR3 orbit, but this star was previously identified as spectroscopic binary (G 127--33, Latham et al., 2002) with a period of 3977 days and it is listed in the SB9 catalog (Pourbaix et al., 2004) with a large RV amplitude of 2.7 km/s. The Gaia astrometric period of 1009 days is very close to the DR3 time range. It is therefore plausible that the Gaia orbit is spurious and the partially-covered binary-star orbit was interpreted as a low-amplitude orbit at $\sim$1/4 of the true period. We therefore argue that Gaia-ASOI-005 should be re-classified as questionable.
@@ 4.3 Brown-dwarf companion candidates
HD 207740 hosts a brown-dwarf companion candidate. Three Keck HIRES spectra of this source are discussed in Rice & Brewer (2020) but no RVs appear to have been published.
The stars HD 206484 and HD 78631 have DR3 solution types of ‘AstroSpectroSB1’ which means that the Gaia RVS also detected a compatible RV variation. The companion mass estimates from binary_masses in these cases are therefore more reliable and straddle the substellar boundary, i.e. these companions are either high-mass brown dwarfs or very-low-mass stars. For HD 78631 we noted a strong discrepancy between our companion-mass estimate ($\sim$123 $M_{\mathrm{Jup}}$, assuming a dark companion) and the one from binary_masses of $\sim$80 $M_{\mathrm{Jup}}$, which was determined using a non-zero flux ratio (Gaia Collaboration et al., 2023b).
@@
GSC 09436-01089 has a solution type of ‘AstroSpectroSB1’ which means that the Gaia RVS also detected a compatible RV variation. Using the primary-mass estimate of ${M_{1}}=1.16\pm 0.12\,M_{\sun}$ from binary_masses and resampling the DR3 orbit parameters while accounting for all covariances we obtain a companion-mass estimate of ${M_{2}}={74.88}^{+9.20}_{-8.07}\,{M_{\mathrm{Jup}}}$. GSC 09436-01089 B is therefore likely a high-mass brown dwarf but it may also be a very-low-mass star.
@@
Finally, we present G 15-6 as the first independently-validated substellar companion identified by our work. The DR3 orbit implies a massive BD companion in a 268 day orbit. This source had already been identified as a possible brown-dwarf host from RV monitoring by Latham et al. (2002) and it is listed in the SB9 catalogue with a period of 266 days. The matching periods and companion-mass estimates leave little doubt that the same orbital motion is detected both by Gaia and the RVs from Latham et al. (2002). Using the primary mass estimate of ${M_{1}}=0.67\,M_{\sun}$ from Latham et al. (2002) with a 10 % uncertainty, and resampling the DR3 orbit parameters while accounting for all covariances we obtain a companion-mass estimate of ${M_{2}}={62.43}^{+6.16}_{-5.44}\,{M_{\mathrm{Jup}}}$. This companion therefore lies firmly in the substellar range and is the first confirmed brown-dwarf companion that was identified with our machine learning approach.
@@ 4.4 Feature Importance Results
In these section, we study the model predictions using SHAP values (Lundberg & Lee, 2017) to identify both the reasons for the models’ predictions on the entire sample population and our candidate pool, but also for individual samples such as the confirmed brown dwarf the models missed.
@@ 4.4.2 Predictions for populations
[…]
@@
Figure 4: SHAP value distribution for the top five features of the top 50 sources identified in each of the 8 configurations (total number of 8$\times$50 sources); colour indicates distribution of the feature; especially smaller values of the variables are associated with higher SHAP values.
Figure 5: SHAP value distribution for the top five features of in a random sample of 50 sources per configuration; colour indicates distribution of the feature; especially larger values of the variables are associated with smaller SHAP values.
Figure 6: SHAP values for mass_function_msun of candidates and random samples (same samples as in Fig. 5) in relation to their mass_function_msun values and in the background a histogram of the value distribution in the non-single source population
@@ 4.4.3 Predictions for individual systems
Investigating individual sources/solutions, we want to highlight three specific samples: The missed confirmed brown dwarf in LHS 1610, the top candidate around HD 40503, and Gaia DR3 246890014559489792 (Gaia object of interest identifier Gaia-ASOI-016), a small mass_function_msun outlier that was not identified as a candidate by our analysis. This can help us understand the decision process of the models, find potential confounders and help interpret the data of individual samples. In addition, decision plots for the missed brown dwarf around HD 89707 are in the Appendix in Figures 14 and 15. It was likely just barely below the detection threshold scoring highly on multiple models.
@@
Finally, Figures 11 and 12 show the decision plots for Gaia DR3 246890014559489792. Looking at the small mass_function_msun value of 1.367e-6 (cf. Figure 3) it would be conceivable that this is a strong candidate. However, as seen in the decision plots this is not the case because the other features, in particular the radial-velocity error, bp_rp, and goodness-of-fit contribute to an inlier classification.
Figure 7: Decision plot showing feature contribution to model output for the random forest models for the brown dwarf in LHS 1610. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 8: Decision plot showing feature contribution to model output for the XGBoost models for the brown dwarf in LHS 1610. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 9: Decision plot showing feature contribution to model output for the random forest models for HD 40503. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 10: Decision plot showing feature contribution to model output for the XGBoost models for HD 40503. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 11: Decision plot showing feature contribution to model output for the random forest models for GaiaDR3 246890014559489792. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 12: Decision plot showing feature contribution to model output for the XGBoost models for GaiaDR3 246890014559489792. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
@@ 5 Discussion
One predictable drawback of the approach in this regard is the tendency of identifying new candidates that mostly lie within the distribution of already-known candidates in terms of their features. For example, the location of our best candidates in the colour-magnitude diagram (Fig. 3 top) largely coincides with the region where the confirmed substellar companions, i.e. the training sample, resides. This is a common drawback of machine learning approaches and could be remedied by providing a richer ‘training‘ dataset that, e.g., includes more confirmed substellar companions around redder sources (bp_rp > 1.5). In the same vein it has to be mentioned that the presented approach is of course greatly impacted by the very limited number of confirmed samples. These aspects will improve with time as the results of various DR3-orbit follow-up studies are being published and certainly with the release of Gaia DR4, when many more orbits of stars with substellar companions will be released. We anticipate that the iterative identification and confirmation of an increasing number of exoplanets with Gaia orbits, covering growing areas in parameter space, will lead to a positive feedback loop with machine learning approaches, which will help to accelerate the pace of discoveries.
@@
Another indicator of such behavior can be seen in Figure 6. Even though mass_function_msun is the strongest feature overall (with small values being associated with a classification as outlier) a small value alone is not sufficient for a classification as outlier as exemplified by Gaia DR3 246890014559489792 and in Figures 11 and 12.
@@ 6 Conclusions
Using machine learning, we identified 22 high-confidence systems with substellar companions, of which five host extrasolar-planet candidates that were previously known. One of the exoplanet candidates (BD+24 4592, Gaia-ASOI-005) is a false-positive because the Gaia orbit does not correspond to the true orbit of the system, which is a known RV binary.
@@
The presence of a brown dwarf with an astrometric mass of ${M_{2}}={62.43}^{+6.16}_{-5.44}\,{M_{\mathrm{Jup}}}$ orbiting our high-confidence candidate G 15-6 was confirmed by literature RV measurements. We propose that the remaining nine high-confidence brown-dwarf companion candidates identified in this study be followed-up preferentially.
With Gaia we anticipate that machine learning analyses of astrometric orbits will grow into becoming powerful complements to the traditional approaches of cross-checking with existing data and literature and following-up with additional observations when pursuing specific scientific goals.
As the Gaia astrometric-orbit catalogue will grow significantly in DR4, machine learning approaches will become more and more important for prioritising follow-up and identifying interesting systems and ‘outliers‘. In the same way, any analysis of the DR4 epoch data (Gaia astrometric timeseries of all sources will be released with DR4888https://www.cosmos.esa.int/web/gaia/release) that has the goal of identifying non-single stars, e.g. through non-linear motions and drifts, single-, or multi-companion orbits, will have to address the problem of false-positives. The holistic machine learning approach that facilitates the use of as much information as available has many advantages in that respect. In particular, once sufficient training data will be available, it is very likely that machine learning algorithms will automatically identify a sizeable fraction of twin binary stars, which are the principal astrophysical false positives that affect Gaia astrometric planet searches.
@@ Appendix A Auxiliary tables
Table 5: Used and computed fields for the gaia_source table. Instead of reproducing the field descriptions from van Leeuwen et al. (2022) here, we included a direct link to the online documentation.
@@
source_id online documentation not used
ra online documentation
dec online documentation
parallax online documentation confounder
parallax_error online documentation
pmra online documentation confounder
pmdec online documentation confounder
visibility_periods_used online documentation
radial_velocity online documentation
radial_velocity_error online documentation
astrometric_chi2_al online documentation
ipd_gof_harmonic_amplitude online documentation
bp_rp online documentation
phot_bp_rp_excess_factor online documentation
ruwe online documentation
phot_g_mean_mag online documentation confounder
phot_rp_mean_mag online documentation
phot_bp_mean_mag online documentation
astrometric_excess_noise_sig online documentation
astrometric_primary_flag online documentation not used
astrometric_excess_noise online documentation
astrometric_n_good_obs_al online documentation
@@
source_id online documentation not used
nss_solution_type online documentation not used
ra online documentation
dec online documentation
parallax online documentation confounder
parallax_error online documentation
pmra online documentation confounder
pmdec online documentation confounder
period online documentation
period_error online documentation
t_periastron online documentation
t_periastron_error online documentation
eccentricity online documentation
eccentricity_error online documentation
astrometric_n_good_obs_al online documentation
obj_func online documentation
goodness_of_fit online documentation
efficiency online documentation
significance online documentation
@@
Figure 14: Decision plot showing feature contribution to model output for the random forest models for the brown dwarf in HD 89707. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.
Figure 15: Decision plot showing feature contribution to model output for the XGBoost models for the brown dwarf in HD 89707. Features are sorted by importance, feature values are shown on the left. A move of the line to the right in the row corresponding to a feature indicates a positive contribution to a classification as an outlier, i.e. exoplanet or brown dwarf, and vice versa.

### v2

@@ 4.2 Exoplanet candidates
The companion of HD 40503 (Gaia-ASOI-002, Gaia Collaboration et al., 2023b; Holl et al., 2023; Marcussen & Albrecht, 2023; Winn, 2022) continues to await full RV confirmation.
We identify HD 128717 (Gaia-ASOI-009) as a strong candidate for hosting a super-Jupiter, motivating prioritised follow-up of this star.
The companion of * 54 Cas (Gaia-ASOI-003) is an exoplanet candidate discussed in Gaia Collaboration et al. (2023b), however, the corresponding orbital solution was retracted as false positive by the Gaia collaboration in May 2024101010https://www.cosmos.esa.int/web/gaia/dr3-known-issues#FalsePositive, thus this is no longer an exoplanet candidate.
The star BD$+$75 510 (HIP 66071, Gaia-ASOI-001, Holl et al., 2023; Gaia Collaboration et al., 2023b) was followed up with high-precision RVs and a sub-Jupiter mass exoplanet was discovered (Sozzetti et al., 2023). However, also this orbital solution was retracted as false positive by the Gaia collaboration in May 2024. The exoplanet discovered by (Sozzetti et al., 2023) does therefore not have a corresponding orbit solution in Gaia DR3 and this is no longer an exoplanet candidate in the context of our study.
The companion of BD$+$24 4592 (Gaia-ASOI-005) appears to be a strong exoplanet candidate in our analysis on the basis of the Gaia DR3 orbit, but this star was previously identified as spectroscopic binary (G 127--33, Latham et al., 2002) with a period of 3977 days and it is listed in the SB9 catalog (Pourbaix et al., 2004) with a large RV amplitude of 2.7 km/s. The Gaia astrometric period of 1009 days is very close to the Gaia DR3 time range. It is therefore plausible that the Gaia orbit is spurious and the partially-covered binary-star orbit was interpreted as a low-amplitude orbit at $\sim$1/4 of the true period. We therefore argue that Gaia-ASOI-005 should no longer be classified as exoplanet candidate.
@@ 4.3 Brown-dwarf companion candidates
We also identified HD 207740 as host of a high-probability brown-dwarf companion candidate. Three Keck HIRES spectra of this source are discussed in Rice & Brewer (2020) but no RVs appear to have been published.
The stars HD 206484 and HD 78631 have Gaia DR3 solution types of ‘AstroSpectroSB1’ which means that the Gaia RVS also detected a compatible RV variation. The companion mass estimates from binary_masses in these cases are therefore more reliable and straddle the substellar boundary, i.e. these companions are either high-mass brown dwarfs or very-low-mass stars. For HD 78631 we noted a strong discrepancy between the companion-mass estimate from binary_masses of $\sim$80 $M_{\mathrm{Jup}}$, which was determined using a non-zero flux ratio (Gaia Collaboration et al., 2023b) and our estimate of $\sim$123 $M_{\mathrm{Jup}}$. The latter was derived under the wrong assumption of a dark companion, and a discrepancy can therefore be expected.
@@
GSC 09436-01089 has a solution type of ‘AstroSpectroSB1’ which means that the Gaia RVS also detected a compatible RV variation. Using the primary-mass estimate of ${M_{1}}=1.16\pm 0.12\,M_{\sun}$ from binary_masses and resampling the Gaia DR3 orbit parameters while accounting for all covariances [^13] The correlation matrix of the Gaia DR3 orbital parameters is available in the archive and should be taken into account when propagating uncertainties to derived quantities, see e.g. Gaia Collaboration et al. (2023b, Appendix D.3). we obtain a companion-mass estimate of ${M_{2}}={74.88}^{+9.20}_{-8.07}\,{M_{\mathrm{Jup}}}$. GSC 09436-01089 B is therefore likely a high-mass brown dwarf but it may also be a very-low-mass star.
@@
Finally, we present G 15-6 as the first independently-validated substellar companion identified by our work. The Gaia DR3 orbit implies a massive BD companion in a 268 day orbit. This source had already been identified as a binary from RV monitoring by Latham et al. (2002) and it is listed in the SB9 catalogue with a period of 266 days. The minimum companion mass determined by Latham et al. (2002) lies in the substellar range but the companion could not be confirmed as a brown dwarf because of the unknown orbital inclination. The Gaia astrometric orbit has now resolved this ambiguity. The matching periods and companion-mass estimates leave little doubt that the same orbital motion was detected independently by both Gaia and in the RVs from Latham et al. (2002). Using the primary mass estimate of ${M_{1}}=0.67\,M_{\sun}$ from Latham et al. (2002) with a 10 % uncertainty, and resampling the Gaia DR3 orbit parameters while accounting for all covariances13 we obtain a companion-mass estimate of ${M_{2}}={62.43}^{+6.16}_{-5.44}\,{M_{\mathrm{Jup}}}$. This companion therefore lies firmly in the substellar range and is the first confirmed brown-dwarf companion that was identified with our machine learning approach.
@@ 4.4 Feature Importance Results
In this section, we study the model predictions using SHAP values (Lundberg & Lee, 2017) to identify both the reasons for the models’ predictions on the entire sample population and our candidate pool, but also for individual samples such as the confirmed brown dwarf the models missed.
@@ 4.4.2 Predictions for populations
In terms of understanding the predictions on the entire population and the candidate pool, Figures 4 and 5 show the SHAP value distributions for the top five features of the candidates and a random sample of sources, respectively. The colour indicates the feature value, e.g. the radial velocity error. Notably, only smaller feature values of the mass function (mass_function_msun) and radial velocity error are strongly associated with higher SHAP values, indicating a greater influence on classifying a sample as an outlier. This trend suggests that certain features are particularly important when their values are small.
@@
Figure 4: SHAP value distribution for the top five features, i.e. the input parameters such as radial velocity error, of the top 50 sources identified in each of the 8 configurations (total number of 8$\times$50 sources); colour indicates distribution of the feature; especially smaller values of the variables are associated with higher SHAP values. The top five features are determined as having the highest average over the entire sample.
Figure 5: SHAP value distribution for the top five features, i.e. the input parameters such as radial velocity error, in a random sample of 50 sources per configuration; colour indicates distribution of the feature; especially larger values of the variables are associated with smaller SHAP values. The top five features are determined as having the highest average over the entire sample.
Figure 6: SHAP values for mass_function_msun of top candidates identified by our approach vs. a random sample of 50 sources per configuration (total number of 8$\times$50 sources) in relation to their mass_function_msun values and in the background a histogram of the value distribution in the non-single source population ; Note that all groups show no trivial relationship between small mass_function_msun values and high SHAP values but form clusters
@@ 4.4.3 Predictions for individual systems
To investigate individual sources/solutions, we want to highlight three specific samples: The missed confirmed brown dwarf in LHS 1610, the top candidate around HD 40503, and Gaia DR3 246890014559489792 (Gaia object of interest identifier Gaia-ASOI-016), a small mass_function_msun outlier that was not identified as a candidate by our analysis. This can help us understand the decision process of the models, find potential confounders and help interpret the data of individual samples. In addition, decision plots for the missed brown dwarf around HD 89707 are in the Appendix in Figures 14 and 15. It was likely just barely below the detection threshold scoring highly on multiple models.
Additionally, Figures 7 - 12 present decision plots that illustrate the contributions of individual features to the model’s prediction. Each row corresponds to a specific feature, and the features are sorted by their overall importance averaged across models. The feature values are displayed on the left-hand side of the plot. For each feature, the coloured line shows how it contributes to the model’s prediction. A positive slope, i.e. an upwards rightward displacement, of the line within a row indicates a positive contribution to the classification of the sample as an outlier like for, e.g., an exoplanet or brown dwarf. Conversely, a negative slope, i.e. and upwards leftward displacement, indicates a negative contribution to the classification. The grey line at the centre of the plot represents the base value of the model’s prediction before considering the effects of individual features. This visualisation helps identifying which features drive the model’s decision and how their specific values influence the output.
@@
Finally, Figures 11 and 12 show the decision plots for Gaia DR3 246890014559489792 / Gaia-ASOI-016. Looking at the small mass_function_msun value of 1.367e-6 (cf. Figure 3) it would be conceivable that this is a strong candidate. However, as seen in the decision plots this is not the case because the other features, in particular the radial-velocity error, bp_rp, and goodness-of-fit contribute to an inlier classification.
Figure 7: Decision plot for the random forest models for the known but not identified brown dwarf in Gaia DR3 43574131143039104 / LHS 1610 (Fitzmaurice et al., 2024); high radial-velocity error, mass_function_msun and bp_rp are the main contributors to the decision against an outliers classification for the models trained on D1, D3 and D4.
Figure 8: Decision plot for the XGBoost models for the known but not identified brown dwarf in Gaia DR3 43574131143039104 / LHS 1610 (cf. Figure 7); high radial-velocity error, mass_function_msun and bp_rp are the main contributors to the decision against an outliers classification for the models trained on D1, D3 and D4.
Figure 9: Decision plot for the random forest models for top candidate Gaia DR3 2884087104955208064 / HD 40503. Small values of the radial-velocity error and mass_function_msun were the most decisive features. Notably, the bp_rp colour was not an important feature.
Figure 10: Decision plot for the XGBoost models for top candidate Gaia DR3 2884087104955208064 / HD 40503. Small values of the radial-velocity error and mass_function_msun were the most decisive features. Notably, the bp_rp colour was not an important feature.
Figure 11: Decision plot for the random forest models for Gaia DR3 246890014559489792 / Gaia-ASOI-016, a candidate with low mass_function_msun. Notably, the other features - especially the radial-velocity error and bp_rp - lead to an inlier classification.
Figure 12: Decision plot for the XGBoost models for Gaia DR3 246890014559489792 / Gaia-ASOI-016, a candidate with low mass_function_msun. Notably, the other features - especially the radial-velocity error, bp_rp, and goodness-of-fit - lead to an inlier classification.
@@ 5 Discussion
One predictable drawback of the approach in this regard is the tendency of identifying new candidates that mostly lie within the distribution of already-known candidates in terms of their features. For example, the location of our best candidates in the colour-magnitude diagram (Fig. 3 top) largely coincides with the region where the confirmed substellar companions, i.e. the training sample, resides. This is a common drawback of machine learning approaches and could be remedied by providing a richer ‘training‘ dataset that, e.g., includes more confirmed substellar companions around redder sources (bp_rp > 1.5). In the same vein it has to be mentioned that the presented approach is of course greatly impacted by the very limited number of confirmed samples. These aspects will improve with time as the results of various Gaia DR3-orbit follow-up studies are being published and certainly with the release of Gaia DR4, when many more orbits of stars with substellar companions will be released. We anticipate that the iterative identification and confirmation of an increasing number of exoplanets with Gaia orbits, covering growing areas in parameter space, will lead to a positive feedback loop with machine learning approaches, which will help to accelerate the pace of discoveries.
@@
Another indicator of such behaviour can be seen in Figure 6. Even though mass_function_msun is the strongest feature overall (with small values being associated with a classification as outlier) a small value alone is not sufficient for a classification as outlier as exemplified by Gaia DR3 246890014559489792 / Gaia-ASOI-016 and in Figures 11 and 12.
@@ 6 Conclusions
Using machine learning, we identified 22 high-confidence systems with substellar companions, of which five host extrasolar-planet candidates that were previously known. Two of those candidates were invalidated during the review process of this paper because the corresponding Gaia orbital solutions were retracted as false positives. We identified one of the exoplanet candidates (BD+24 4592, Gaia-ASOI-005) as a false-positive because the Gaia orbit does not correspond to the true orbit of the system, which is a known RV binary.
@@
The presence of a brown dwarf with an astrometric mass of ${M_{2}}={62.43}^{+6.16}_{-5.44}\,{M_{\mathrm{Jup}}}$ orbiting our high-confidence candidate G 15-6 was confirmed by literature RV measurements. We propose that the remaining nine high-confidence brown-dwarf companion candidates identified in this study be followed-up preferentially, e.g. with radial velocities or interferometric imaging (e.g. Winterhalder et al., 2024).
As the Gaia astrometric-orbit catalogue will grow significantly in DR4, machine learning approaches will become more and more important for prioritising follow-up and identifying interesting systems and ‘outliers‘. In the same way, any analysis of the DR4 epoch data (Gaia astrometric timeseries of all sources will be released with DR4151515https://www.cosmos.esa.int/web/gaia/release) that has the goal of identifying non-single stars, e.g. through non-linear motions and drifts, single-, or multi-companion orbits, will have to address the problem of false-positives. The holistic machine learning approach that facilitates the use of as much information as available has many advantages in that respect. In particular, once sufficient training data will be available, it is very likely that machine learning algorithms will automatically identify a sizeable fraction of twin binary stars, which are the principal astrophysical false positives that affect Gaia astrometric planet searches.
@@ Acknowledgements
The code and data needed to reproduce our results are available at https://github.com/esa/gaia-astrometric-exoplanet-orbit-ml.
@@ Appendix A Auxiliary tables
Table 5: Used and computed fields for the gaia_source table. Instead of reproducing the field descriptions from van Leeuwen et al. (2022) here, we included direct links to the online documentation. Columns commented as ‘not used‘ and ‘confounder‘ were not used in the analyses. The latter correspond to features that are marked as important for identifying outliers but that are not expected as such based on physical arguments, see Section 3.3.
@@
source_id Identifier not used
ra Right Ascension (RA)
dec Declination (Dec)
parallax Trigonometric parallax confounder
parallax_error Standard error
pmra Proper motion in RA confounder
pmdec Proper motion in Dec confounder
visibility_periods_used Number of independent measurements
radial_velocity Radial velocity
radial_velocity_error Standard error
astrometric_chi2_al Fit-quality metric
ipd_gof_harmonic_amplitude Statistic related to scan-angle dependence
bp_rp Stellar colour
phot_bp_rp_excess_factor Colour excess factor
ruwe Renormalised fit-quality
phot_g_mean_mag Apparent magnitude in $G$ confounder
phot_rp_mean_mag Red Gaia magnitude ()
phot_bp_mean_mag Blue Gaia magnitude ()
astrometric_excess_noise Source-specific excess noise
astrometric_excess_noise_sig Excess noise significance
astrometric_primary_flag Primary source flag not used
astrometric_n_good_obs_al Number of used measurements
@@
source_id Identifier not used
nss_solution_type Solution type not used
ra RA
dec Dec
parallax Trigonometric parallax confounder
parallax_error Standard error
pmra Proper motion in RA confounder
pmdec Proper motion in Dec confounder
period Orbital period
period_error Standard error
t_periastron Periastron time
t_periastron_error Standard error
eccentricity Orbit eccentricity
eccentricity_error Standard error
astrometric_n_good_obs_al Number of used measurements
obj_func $\chi^{2}$ metric
goodness_of_fit Gaussianised $\chi^{2}$ metric
efficiency Correlation metric
significance Orbit significance $=a_{0}/\sigma_{a_{0}}$
@@
Figure 14: Decision plot showing feature contribution to model output for the random forest models for the brown dwarf in HD 89707. Radial-velocity error and mass_function_msun were the most decisive features where the most important features.
Figure 15: Decision plot showing feature contribution to model output for the XGBoost models for the brown dwarf in HD 89707. Radial-velocity error and mass_function_msun were the most decisive features where the most important features.